OpenStudy (anonymous):

Ok, back again for more calc and derivatives. Question and my answer will be in the comments. (Thanks to everyone that spends time helping me and others. You are saints among men!)

OpenStudy (anonymous):

so the question is\[\frac{ e ^{x}-1 }{ \sqrt[3]{x ^{2}+x} }\] and my answer is \[\frac{ e ^{x} }{ (x ^{2}+x)^{1/3} } + \frac{ e ^{x}-1 }{ (-\frac{ 1 }{ 3 }(x ^{2}+x)^{2/3} }\]

OpenStudy (anonymous):

that does not look like a question to me. What are we supposed to do? :/

OpenStudy (anonymous):

I am wondering if i got the correct answer. The first equation is the original problem, and the second equation is my answer f'(x)=... I did it using the product rule

OpenStudy (anonymous):

\[\frac{ e^{x}-1 }{ \left( x^{3}+x \right)^{\frac{ 1 }{ 3 }} } - \frac{ (e ^{x}-1)(3x^2+1) }{ 3(x^3 +x)^\frac{ 4 }{ 3 } }\]

OpenStudy (anonymous):

this is what I get with the product rule

OpenStudy (anonymous):

why not use the quotient rule?